OpenStudy (anonymous):

The area of a triangle is 67. If the height was 13, find the value of the base. Answer 20.62 9.89 10.31 2.58

OpenStudy (anonymous):

A=(1/2)B*H 67=(1/2)B*(13) (67*2)/(13)=B B=10.307 rounded would give you 10.31

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

no problem